### v1

Abstract
Scheduling virtual machines (VMs) to hosts in cloud data centers dictates efficiency and is an NP-hard problem with incomplete information. Prior work improved VM scheduling with predicted VM lifetimes. Our work further improves lifetime-aware scheduling using repredictions with lifetime distributions vs. one-shot prediction. The approach repredicts and adjusts VM and host lifetimes when incorrect predictions emerge. We also present novel approaches for defragmentation and regular system maintenance, which are essential to our data center reliability and optimizations, and are unexplored in prior work. We show that repredictions deliver a fundamental advance in effectiveness over one-shot prediction.
We call our novel combination of distribution-based lifetime predictions and scheduling algorithms Lifetime Aware VM Allocation (LAVA). LAVA improves resource stranding and the number of empty hosts, which are critical for large VM scheduling, cloud system updates, and reducing dynamic energy consumption. Our approach runs in production within Google’s hyperscale cloud data centers, where it improves efficiency by decreasing stranded compute and memory resources by $\sim$3and $\sim$2respectively, and increases availability for large VMs and cloud system updates by increasing empty hosts by 2.3-9.2 pp in production. We also show a reduction in VM migrations for host defragmentation and maintenance. In addition to our fleet-wide production deployment, we perform simulation studies to characterize the design space and show that our algorithm significantly outperforms the state of the art lifetime-based scheduling approach.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Cloud data centers run an increasing fraction of the world’s compute. Efficient resource use in data centers is thus critical, from both an economic and environmental perspective. In these Infrastructure-as-a-Service (IaaS) environments, VM scheduling to physical hosts determines a large portion of end-to-end efficiency. Poor assignment of VMs may strand host resources – remaining host resources are too small or too imbalanced to accommodate additional VMs. VM allocation also impacts the number of empty hosts, which impacts idle power optimizations, resource utilization, the ability to schedule large VMs (obtainability), and system maintenance (e.g., for security patches).
Much prior work addresses workload allocation in data centers Schwarzkopf et al. (2013); Verma et al. (2015); Cortez et al. (2017); Gog et al. (2016); Grandl et al. (2016); Zhang et al. (2014), including leveraging workload predictions Delimitrou & Kozyrakis (2014b; 2013); Park et al. (2018). The closest related work predicts job lifetimes (time between start and exit) to assist the scheduler in making better placement decisions Yadwadkar et al. (2014); Ferguson et al. (2012); Tumanov et al. (2016); Cortez et al. (2017); Buchbinder et al. (2021). Most recently, Barbalho et al. Barbalho et al. (2023) show bin packing improvements using Lifetime Alignment (LA) scheduling. LA predicts the lifetime of a VM at creation time, and then treats the predicted lifetime as fixed. As a result, mispredictions may tie up an entire host and, over time, model mistakes accumulate. For example, assume a model with a 1% chance of predicting a long-lived VM (running for months) as short-lived. If, over time, the scheduler places 70 presumably short-lived VMs on a host, there is a 50% chance that one of them is long-lived and prevents the host from freeing up until that VM exits. Correspondingly, Barbalho et al. show that their best algorithm achieves maximum efficiency with perfect prediction accuracy and degrades rapidly as the accuracy decreases (Barbalho et al. (2023), Figure 8). They thus introduce an alternative algorithm that achieves lower efficiency but is more tolerant to mispredictions.
Our work achieves both tolerance to mispredictions and high efficiency, by repredicting VM and host lifetimes to correct mispredictions and improve accuracy over time. We take inspiration from work in C++ memory allocation that increases robustness of lifetime-based allocation by reassessing predictions on the fly Maas et al. (2020). In addition to bin packing, we address significant challenges not explored by prior work, including host defragmentation, maintenance, and stranding, which are essential components of Google’s cloud system design and approach to efficiency.
Our key contributions are A) VM lifetime reprediction and adaptation to mispredictions, such as when a VM outlives its initial prediction; B) a novel ML model of VM lifetime probability distributions, which captures hard-to-predict VM behavior and prediction uncertainty; and C) three novel lifetime repredicting scheduling algorithms: 1) Lifetime Aware Scoring (LAS) incorporates predictions into an existing scoring function and corrects mispredictions. 2) Lifetime Aware VM Allocation (LAVA) more fundamentally redesigns the scheduler around lifetimes. 3) Lifetime-Aware ReScheduling (LARS) exploits lifetimes to reduce VM disruptions during host defragmentation and maintenance.
Our approach is deployed fleet-wide across a planet-scale cloud infrastructure at Google. We evaluate our approach in production and through simulations with production traces, and compare to prior work. We validate that our simulations have high fidelity by comparing against production behavior.
  • This paper shows that our key insight – probability distribution models of VM lifetimes – creates opportunities for new co-designed algorithms that repredict VM lifetimes and actively respond to mispredictions.
  • It shows the new LAS, LAVA, and LARS algorithms increase empty host availability, reduce stranding, and reduce VM disruptions.
  • Using production data and extensive simulation studies, we show that these algorithms are efficient and robust, outperforming the state of the art (SOTA).
  • We incorporate lifetime-based scheduling in a complex environment that features dynamic resource management based on VM usage, VM live migrations, and hierarchical scheduling. We show reductions in resource stranding and VM migrations, in addition to improved bin packing quality.
Similar to Maas et al. Maas et al. (2020), we find on-the-fly adjustment to mispredictions, versus one-off predictions, opens up a fundamentally different class of algorithms and optimizations, in this case for VM scheduling, and perhaps beyond. Our production deployment shows the practicality of this approach, and of deploying ML in the lower layers of the systems infrastructure stack more broadly.
2 Background
This section overviews VM lifetime properties, our production system, and metrics. Next, it describes the SOTA Lifetime Alignment algorithm (LA) Barbalho et al. (2023), to which we compare qualitatively and quantitatively.
2.1 Characterizing VM Lifetimes in Production
In our production environment, most VMs are short-lived, but most core hours are consumed by long-running VMs, similar to the generational hypothesis in garbage collection Lieberman & Hewitt (1983); Ungar (1984). Figure 1 shows our VM lifetime distributions. While 88% of all VMs live for less than 1 hour, 98% of resources are consumed by VMs that live for 1 hour or more (measured in CPU cores $\times$ time occupied). Without lifetime-specific handling of VMs, each VM would introduce the same amount of fragmentation, although only the placement of a small fraction of them has an impact on the overall resource efficiency. Prioritizing the bin packing quality of all VMs equally is thus highly inefficient, and motivates lifetime-based VM scheduling.
A key element of our approach is to view VM lifetimes from the perspective of distributions (Cumulative Density Functions or CDFs), as opposed to averages. We identify that VM features useful for prediction do not fully determine VM lifetimes. The best a traditional model could do would thus be to predict the expected (or average) lifetime for each VM. Figure 2 shows the average lifetime may not be particularly meaningful when there are both long and short-lived VMs with the same features. We therefore model and predict the lifetime of VMs as probability distributions. Note that this type of approach has a long history in survival analysis Wang et al. (2019) and in areas such as storage systems Zhou & Maas (2021). However, our novelty is predicting distributions and co-designing VM scheduling algorithms to leverage them in detecting and updating mispredictions.
Figure 1: Distribution of VM lifetimes of scheduled VMs vs. their resource consumption.
Figure 2: Lifetime distribution (PDF) of VM lifetimes. When the VM is scheduled, the expected (average) lifetime is 0.2 days. After it has run for 1 day, the expected remaining lifetime is 4 days. After 7 days, the expected remaining lifetime is 10 days.
2.2 VM Allocation
Our fleet consists of data centers in many geographic regions that run a mix of first and third party workloads. Cloud workloads have distinct host pools for distinct VM families. VM families represent different products, such as VMs optimized for high performance or cost efficiency. Broadly, VM families fall into two categories. 1) Slice-of-hardware families assign each VM a fixed partition of a host’s resources, e.g., CPU cores, DRAM, SSD, and GPUs. 2) Dynamically sized families share unused resources based on VM and host usage patterns and thus the amount of virtual CPUs and memory of VMs may exceed the physical resources on a host. This paper focuses on two VM families, C2 and E2, representative of these categories.
A scheduling framework called Borg Verma et al. (2015) assigns VMs to hosts within pools. Borg maintains a global view of all VMs and hosts in a given pool. When a VM creation request arrives, the scheduler computes the set of feasible hosts, i.e., hosts that have sufficient resources and match any hard constraints. For each feasible host, it then computes a score that determines the preference between them. This score incorporates a number of different business goals, such as spreading tasks across failure domains and bin packing efficiency. This scheduling setup is common Barbalho et al. (2023); Hindman et al. (2011); Verma et al. (2015).
Borg’s Waste Minimization bin packing algorithm explicitly optimizes for producing empty shapes of resources in multiple dimensions (e.g., CPU $\times$ DRAM $\times$ SSD) that match anticipated workload patterns. This approach was deployed several years ago based on experiments showing it achieves higher obtainability for large VMs and improves resource efficiency compared to Best Fit, the previous scheduler, which was similar to prior work Barbalho et al. (2023). However, Borg’s scoring function poses challenges for incorporating VM lifetimes, because it uses lexicographic ordering – one scoring dimension is evaluated at a time, with tie-breakers resolved by the next-lower scoring function. Bin packing is among its final dimensions, ensuring that improving bin packing does not degrade other business objectives.
2.3 Constraints & Objectives
Our work aims to maximize resource utilization and minimize fragmentation without degrading other business objectives. We measure the following optimization metrics.
Empty Hosts. We measure the number of empty hosts in a pool as a percentage of the total number of hosts. Empty hosts are an important resource as they are required to schedule large VMs that consume most or all of the host’s resources. They also enable host maintenance (e.g., rolling out a kernel or microcode security patch) without disrupting VMs with live migrations. Additionally, Google may put empty hosts in low power mode or move empty hosts between pools and configure them differently to serve various capacity needs. As such, 1 pp (percentage point) of increase in empty hosts directly corresponds to 1% more capacity available for these use cases. Other metrics of bin packing quality include packing density Barbalho et al. (2023), and are equivalent.
Resource Stranding. This metric measures free resource shapes and how many future VMs may be scheduled in them. Periodic inflation simulations measure stranding. We take a representative mix of VMs and simulate scheduling as many of them as possible until capacity is exhausted. The remaining resources on hosts represent stranded resources that cannot fit new VMs. For example, a host may contain free memory but no free CPUs. This memory is thus stranded. The metric is similar to admission control scoring functions Sajal et al. (2023) and captures unusable capacity.
VM Live Migrations. We aim to minimize live migrations. Live migration and fragmentation are closely connected. For some VM families, periodic live migrations will defragment resources. Live migration and resource efficiency may thus be traded against one another. This trade-off is controlled by policy and predates our work. It represents a fundamental difference from some other clouds. Maintenance events often trigger live migrations as well.
2.4 Baseline: Lifetime Alignment (LA)
We compare against LA (the Lifetime Alignment algorithm) from Barbalho et al. Barbalho et al. (2023). While other work on lifetime-based VM scheduling predates it, LA supersedes it and, to our knowledge, is the only published approach deployed in production data centers. At a high level, our setup (Figure 3) resembles Barbalho et al. We use ML to predict the lifetimes of VMs and incorporate predictions into scheduling decisions. Note that they use Best Fit scoring and do not consider maintenance and defragmentation cases, contributions of our paper.
For every VM allocation request, LA runs a gradient-boosting model to predict its lifetime class, which is over an exponentially increasing time interval ($2^{i}$ to $2^{i+1}$). Their model run on special inference servers. The allocator then preferentially assigns VMs to a host with the same lifetime class, using Best Fit. The lifetime class of a host is defined dynamically: it is the longest remaining time of any VM on that host, based on the VM’s initial lifetime prediction. If there is no host of the same lifetime class, LA picks another suitable host. If no such host exists, it assigns the VM to a new (previously empty) host.
Barbalho et al. report that LA achieves high improvements in packing density with accurate predictions. While LA performs live migrations occasionally, they are not reported or part of the scheduling approach Barbalho et al. (2023). Because they find the multi-class version of LA is sensitive to mispredictions, they introduce DPBFR, which uses lifetime predictions only to adjust the quantization of Best Fit scoring. They deployed DPBFR in production, as it is more robust to mispredictions, but it achieves lower maximum improvement compared to the binary version of LA. We thus compare to a faithful implementation of their best algorithm, called LA-Binary.
Figure 3: Overview of our VM allocation setup.
While LA Barbalho et al. (2023) tackled optimizing packing density (empty hosts), we extend the state-of-the-art by reducing stranding and our LARS algorithm decreases the number of migrations needed due to maintenance and defragmentation by determining the migration order. Since we create more empty hosts via lifetime-aware scheduling, we reduce the need for defragmentation.
2.5 Lifetime-aware C++ Memory Allocation
We observe that VM scheduling is similar to explicit memory allocation in C/C++ and other programming languages. The memory allocator needs to place objects, that never move, on pages such that pages are either maximally used or become fully empty Maas et al. (2020). VM scheduling is similar. VMs correspond to objects, hosts to pages, stranding to internal fragmentation, and empty hosts to external fragmentation. A key difference is that VM allocation is multi-dimensional. It manages multiple resources, including CPUs, memory, and SSD, while memory allocation is one-dimensional (object size).
LLAMA Maas et al. (2020) is a C++ memory allocator that uses an ML model to predict object lifetimes. Every page is assigned a lifetime class (10ms, 100ms,…) and objects are initially assigned to huge pages with the same lifetime class. Once a page is full, emerging gaps are filled with objects that are at least one lifetime class lower. Once all original objects are free, LLAMA reduces the lifetime class of the page by one. This process continues until the page is empty. LLAMA detects mispredictions. If a page is not free before the expected time, LLAMA under-predicted the lifetime of an object, and it increases the lifetime class of the page by one to correct the misprediction. This mechanism minimizes fragmentation while tolerating mispredictions. The latter is critical, since a single misprediction may fragment an entire huge page and misprediction probability increases exponentially with the number of allocations.
We take inspiration from LLAMA, designing our models and allocators to detect and correct mispredictions. We cannot directly adopt LLAMA because VM allocation is multi-dimensional and bin packing is not the only business objective. We thus incorporate the key insights of LLAMA into new algorithms, LAS, LAVA, and LARS.
3 ML Model
We design ML models that predict the remaining lifetime of a VM both at the time it is scheduled and when re-evaluating the VM and host later, to correct previous mispredictions. Intuitively, reprediction addresses two fundamental issues.
  1. 1.
    VM lifetimes depend on many factors that cannot be described by its features (e.g., human behavior or external events). As such, some VMs are fundamentally impossible to predict. Predicting a single VM lifetime thus limits accuracy as shown in Figure 2.
  2. 2.
    Estimating the actual lifetime of a previously scheduled VM throughout its execution is not always possible with a single prediction. For example, if a particular VM type’s lifetime is bi-modal (e.g., 1 day or 1 week), observing that the VM has been running for 2 days means that it is expected to live for another 5 days.
Our model addresses these limitations by predicting a probability distribution as opposed to a single value. This approach is well-known in survival analysis Wang et al. (2019). Survival models capture the “uncertainty” associated with a particular set of feature values. 1) If a VM is highly predictable based on the features, the distribution will be narrow, while difficult-to-predict VMs have wider distributions. 2) We use distributions to update our VM estimates based on the lifetime so far (uptime). Instead of relying on the initial prediction, we compute the conditional expectation $\mathbb{E}(T_{r}|T_{u})$ for an updated estimate of a VM’s lifetime over time: “Given a VM has been running for interval $T_{u}$, what is the expected remaining lifetime $T_{r}$?”. This value can be directly calculated from the PDF in Figure 2.
This approach avails itself to a range of models, including different forms of survival models, neural networks, and gradient-boosted decision trees (GBDTs). GBDTs performed best in our experiments. Since they are not distribution-based, we pass an additional parameter $T_{u}$ into the model that represents the uptime of the VM so far (i.e., the x axis of the PDF, where lifetime = uptime + remaining lifetime), and predict the remaining lifetime. We augment every training example by turning it into multiple examples with different values of $T_{u}$: specifically, 12.5%, 25%, … of the original lifetime. This implementation turns a regression model into a survival model with good accuracy. We train the model directly on $\mathbb{E}(T_{r}|T_{u})$ as the label.
There are two ways to use such a model: 1) When scheduling a VM, run the model repeatedly to materialize a mapping from uptime to the expected remaining lifetime, which may require a large number of model invocations to get sufficient precision. 2) Run the model repeatedly when re-evaluating a VM, which saves resources since it only runs the model for values that are actually needed, but only works if model latency is low. We choose the latter.
Our models use a range of features, including the zone, VM family, VM shape, whether or not the VM attaches to local SSD, and internal metadata. We use a production-grade machine learning library called Yggdrasil Guillame-Bert et al. (2023). We generate our training data by querying a large-scale internal data warehouse that contains historical data of VMs. We use a combination of distributed SQL queries and data processing pipelines to generate a set of labeled examples and feed them into the training framework. Our data sets are on the order of 1M examples for training. The resulting model is loaded from a file system and runs in binaries via a range of language bindings. We periodically retrain our model on recent data.
We train a single joint model for our entire global fleet, across a large number of our data centers. We experimented with training a separate model for each pool, but this approach is disadvantageous from a rollout perspective, more difficult to maintain, and we found that it neither performs better nor significantly improves resource savings. Our model achieves 99% precision at 70% recall when used to classify VMs between short and long-lived according to a 7 day threshold. We use this model throughout all of our experiments, to ensure comparability across all algorithms.
An important design consideration is how to deploy the models. For instance, Barbalho et al. Barbalho et al. (2023) run their models on separate inference servers. They cache prediction results to tolerate unavailability of servers, increasing complexity. Running model inference on separate servers also requires handling model rollouts and verification separately since they are not integrated with existing rollout mechanisms.
We choose instead to embed the model directly into the Borg binary. This deployment reduces possible failure modes. By piggybacking on Borg, we update models with the same frequency as Borg, testing the two together. This approach lowers model overhead and achieves a median latency of 9 us (Figure 8), which is 780$\times$ lower than the median latency for LA Barbalho et al. (2023). With this low latency, we can run our model more frequently, including for lifetime-guided defragmentation (Section 4.4) and to update and correct mispredictions.
Our model has been deployed in production for over 9 months and has maintained all our production standards, while improving empty machines, reducing stranding, and reducing VM migrations.
4 Scheduling Algorithms
We show that repredicting lifetimes offers a fundamental advantage over one-shot prediction in our multi-dimensional scheduling approach. We introduce three algorithms. LAS is non-invasive and becomes active when a set of hosts satisfy all other criteria and thus have the same score without the bin packing score (Section 4.2). LAVA, in contrast, makes lifetime-aware scheduling a fundamental part of the scheduler (Section 4.3). LARS exploits lifetimes for defragmentation and maintenance (Section 4.4).
4.1 Theoretical Foundation
We can show analytically that repredicting lifetimes and correcting for mispredictions fundamentally improves how well any given algorithm can do. Specifically, if the initial error in lifetime prediction is a positive constant, then the number of hosts required without correcting for mispredictions will exceed the number for the same best fit algorithm with repredictions by $\Omega(m)$, where $m$ is the number of hosts.
4.2 LAS: Non-Invasive Lifetime-Aware Scheduling
The goal of LAS (Figure 4) is to create more empty machines by scheduling VMs on a host where all other VMs are likely to exit at a similar time or later. Given a VM, LAS ranks potential hosts based on the exit times for the VMs already scheduled on that host. It computes the maximum of the predicted exit time of all VMs on the host. This score is conceptually similar to LA Barbalho et al. (2023), with a crucial difference. Instead of using the original VM lifetime predictions, LAS repredicts the remaining lifetime (i.e., the updated exit time) for all VMs, using their uptime so far.
LAS corrects any earlier mispredictions as follows. Consider a host where all remaining VMs have exceeded their originally predicted lifetime. Without correcting mispredictions, the scheduler has to assume that this host will soon become empty and will avoid scheduling long-lived VMs on this host. We hypothesize that this choice is the reason why LA is sensitive to prediction accuracy, leading the authors to not deploy LA. In contrast, LAS schedules new VMs on hosts using the repredicted remaining lifetimes of all VMs on each host. As such, LAS continues to improve its predictions if the original prediction was wrong.
Figure 4: Overview of the LAS algorithm.
[…]
We add the temporal cost into our lexicographic scoring function one level above the bin packing score. The temporal cost never has an effect on any higher-ranked business metrics. It only becomes the deciding factor if all those scores are the same. The score is thus non-invasive. It is entirely local to the bin packing score and is equivalent to modifying the bin packing score to take a third dimension (lifetime) into account.
4.3 LAVA
A limitation of LAS is that mispredictions may gradually bump up the host lifetime. By trying to match the new VM to the exit times of other VMs on a host, the host may never free up because each under-predicted VM further bumps up the exit time, causing even more VMs to be scheduled, which in turn bumps up the exit time further.
In preference to filling resource gaps on hosts with VMs that exit at a similar time, LAVA attempts to fill gaps with VMs that are at least 10$\times$ shorter lived, so that even mispredictions are unlikely to increase the host lifetime. Further, LAVA updates the host’s exit time when detecting a major (more than 10$\times$) misprediction. We use order of magnitude differences to limit the impact of small amounts of over or under prediction. We now describe LAVA in more detail.
LAVA divides lifetime predictions into lifetime classes: $<$1h, 1-10h, 10-100h, 100-1000h. We refer to these lifetime classes as LC1, LC2, L3, and L4. Each host is also assigned a lifetime class. LAVA runs an additional coarse-grained host scoring function based on their lifetime classes and breaks ties with LAS.
Each host has a lifetime class and is in one of three states: empty, open, or recycling (similar to LLAMA Maas et al. (2020)). When a new VM arrives and there is no eligible open or recycling host, LAVA places the VM on an empty host, assigns the host the open state, and assigns the host the lifetime class of the VM. As long as this host is open, LAVA will only schedule other VMs of the same lifetime class on it (Figure 5(a)). Once over 90% of the resources (CPU or memory) of an open host are occupied, LAVA transitions the host to the recycling state. All the VMs that were present on the host when it transitioned state are labeled residual VMs (i.e., the VMs of the original lifetime class of the host). Recycling hosts only receive VMs that are at least one lifetime class lower, e.g., LAVA only places new VMs with a predicted lifetime of 10 hours or shorter on 100-hour recycling hosts.
LAVA handles mispredictions as follows. If all predictions are correct, the total lifetime of a host does not exceed 1.1$\times$ its original lifetime class. There are two cases: overpredicted and underpredicted lifetimes.
If lifetimes are overpredicted, VMs will exit early, creating new gaps on the host. The approach of placing shorter-lived VMs in gaps ensures that we keep a host occupied, while minimizing the likelihood that we extend its lifetime. Once all residual VMs exit, we know that all remaining VMs are of the next-shorter lifetime class. We therefore re-classify the host as one lifetime class lower (Figure 5(b)), classify all remaining VMs as the new residual VMs, and fill emerging gaps with even shorter-lived VMs. This process repeats until the host is empty.
Conversely, when lifetimes are underpredicted, the host will not become empty within the expected time interval. We detect this case by establishing a time-out for each host after which we move it to the next-higher lifetime class (Figure 5(c)). All VMs that are on the host at the time of expiration become the new residual VMs.
Building upon LAS, LAVA is integrated into the existing hierarchical scheduler by adding another scoring function before LAS. This higher-ranked scoring function keeps track of additional states such as predicted host lifetime class, host state, and deadline for each host. The size of this additional state is negligible compared to the substantial amount of existing per-VM state.
When a new VM request arrives, LAVA scores the hosts in the following order (in decreasing preference): recycling hosts whose lifetime classes are greater than the VM lifetime class (the closer the more preferred), open hosts whose lifetime class are equal to the VM lifetime class, any non-empty host, and lastly empty hosts. In contrast to LAS, which acts as a tie-breaker for the existing algorithm, LAVA makes decisions that the baseline could not have made, and is therefore invasive.
(a) LAVA preferentially places new VMs on recycling hosts of a higher lifetime class. If none is available, it chooses open hosts of the same lifetime class. Otherwise, it opens a new host.
(b) When all residual VMs (highlighted) on a host exit, LAVA reduces the lifetime class by one. All remaining VMs turn residual.
(c) If a host exceeds its deadline, there was a misprediction and LAVA increases its lifetime class by one. Its VMs become residual.
Figure 5: Overview the LAVA Algorithm.
4.4 LARS: Defragmentation & Maintenance
The above algorithms focus on reducing fragmentation by improving VM placement. Borg also performs software and hardware maintenance activities, applying security, microcode, host OS, and VM software updates. It also actively defragments hosts for some VM families. These activities live migrate VMs. Live migration uses pre- and post-copy techniques that minimize VM execution time disruptions to seconds. However, total migration times are proportional to VM memory and SSD usage, and consume capacity on both hosts. In simulations, we model that both hosts are busy for a conservative 20 minutes.
Borg reserves a small fraction of hosts for defragmentation, maintenance, and to tolerate hardware failures, and thus these activities do not impact available capacity. We show how using lifetimes can reduce the number of VMs that are disrupted during defragmentation and maintenance. We call the algorithm LARS. The rest of our exposition focuses on defragmentation for conciseness.
For some VM families, we defragment hosts when the availability of empty hosts in a particular pool drops below a certain threshold. The defragmenter searches for suitable hosts to defragment, preferring hosts with few VMs and excess resources and to move the VMs to hosts with available slots that match the VM shapes.
LARS first chooses a small number of hosts to defragment based on their excess resources and marks them as unavailable for scheduling. It then uses predicted lifetimes to determine the order in which to migrate VMs. It starts with the longest-predicted VM and selects a target host to which to migrate it. Choosing this VM reduces the number of migrations, as shorter-lived VMs exit while LARS migrates longer-lived VMs. Once the host is empty, LARS marks it as available to schedule for any lifetime class.
Note that LARS uses the same algorithm to choose a target host for migration as it does during initial VM placement. As such, LAS and LAVA use repredictions from our distribution-based model to place the VM on a better-suited host than it was placed initially during migration.
5 Implementation
Figure 6: Empty host improvements comparing LAS, LAVA with LA-Binary, to our production baseline for 24 C2 pools in simulation. On average, LAVA increases empty hosts by 6.5 pp, compared to 6.1 pp for LAS, and 5.0 pp for LA-Binary.
Our algorithms are integrated into Borg’s main scheduler code base. The scheduler runs within Borg Prime, which handles incoming VM requests and assigns them to hosts. The Borg Prime binary is replicated across a number of instances, with one of them serving as the elected leader. All replicas run the same binary and consensus is handled at a lower layer. From the perspective of our algorithms, we treat this scenarios as if there was a single Borg Prime instance.
As mentioned in Section 3, our model is compiled into the Borg Prime binary and rolled out with Borg. This coupling has the advantage that canary deployments that are part of our rollout process test the entire end-to-end approach. We deploy our software regularly, and may support updating our model as often as once a week. However, we found that the models’ accuracy remains high even a month later as shown in Figure 10. Note that while we found it highly advantageous from a latency and complexity perspective to compile the model into the binary, our approach is not restricted to this deployment scenario.
We carefully measured the CPU and DRAM overheads of our approach in production. Binary size increases by less than 10 MB, and both the memory and CPU consumption of our approach are indistinguishable from noise. This lack of observable overhead may seem counter-intuitive, but a VM record has so much state and Borg performs so much computation during scoring that the extra bytes of memory and the 9 us of model latency are not visible. We process on the order of 10-100 scheduling requests per second in each cluster, and even hypothetically repredicting 100 VMs in the process would add less than 1 ms. In practice, we repredict much fewer VMs, since most hosts are filtered out by higher-ranking scoring functions and only VMs on considered hosts are re-scored. The computation is parallelized as well.
5.1 Simulations
LAS/LAVA: We developed an event-driven simulator to explore our approach. We extract production traces of VM start, exit, and restart events within a particular pool and then replay this trace against a simulated instance of the scheduler. The simulated scheduler uses our production code base and is thus highly accurate. We validated our simulator against production.
LARS: To calculate the benefit of LARS (lifetime-based defragmentation), we built a modified simulator, using the same production traces. From our traces, we collect the full list of migrations that were performed during a time interval. In production, we limit the number of live migrations that can be in progress at a time to batches of 3. We simulate this behavior by assuming that all migrations are performed in a certain order (in our baseline, defined by the trace), but have to wait until a slot is available. This approach has the effect that some VMs exit while others are migrating. LARS modifies this order based on lifetime predictions.
5.2 Production Measurements
Measuring the impact of scheduling policies in production is challenging. Since workloads and total capacity in a pool are constantly changing, it is difficult to attribute metric changes to a particular modification. We therefore create an A/B testing methodology. We split the hosts into two halves and apply our scheduling algorithm to one of them. Our production metrics report empty hosts, stranding, and model prediction for the entire pool. However, stranding is only accurate for entire pools and cannot differentiate A/B hosts. We report live production experiments in four pools, where we apply the algorithm to the entire pool and evaluate scheduling latency. These experiments also ensure that our A/B approach does not distort behavior. Note that these A/B deployments are at production scale. The cluster/zone sizes for this product range from 100s to 10,000+ hosts with O(100-10,000) VMs active at a time.
LAS has been deployed fleet-wide for one VM family for months, at the time of writing. Time series data from the fleet-wide production rollout confirmed our simulation numbers and A/B experiments, and show a fleet-wide increase of empty hosts by 4 pp for this VM family.
5.3 LA-Binary Comparison
Our main comparison is to LA-Binary Barbalho et al. (2023). Since our data center environment is different, perfectly replicating their setup is impossible, e.g., the ML model features are different. We are faithful to LA-Binary’s description and favorable to LA-Binary when in doubt. For an apples-to-apples comparison, we use the exact same model when comparing LA-Binary to LAS and LAVA, but without repredictions for LA-Binary. VMs are classified as short or long lived based on their lifetime predictions with the same two-hour cutoff threshold as in their paper.
6 Evaluation
This section reports results from production and simulations on production data center traces.
6.1 End-to-end Simulation Results (LAS & LAVA)
We first compare performance of the different algorithms in simulation. These experiments are based on the C2 family. Our simulations are highly accurate.
Figure 6 shows simulation data from C2 production traces from 24 pools that cover a wide range of sizes, geographies, and usage patterns. The traces are from May-June 2024 and have a duration of over seven weeks. We measure the average percentage of empty hosts and report changes relative to our current production baseline. For instance, if LAVA achieves 21% and the baseline achieved 20%, we report a 1 percentage point (pp) improvement. We run all algorithms both with our actual lifetime model and a perfect (oracular) predictor as a comparison point. Similar to the LA paper (which increased packing density by 2% in production), a consistent 1 pp improvement represents a large gain in this scenario, and can be equivalent to 1% of a cluster’s capacity if used to power down or divest hosts, or offset holdbacks.
LAS outperforms LA-Binary on average, achieving an additional 1.1 pp increase in empty hosts. With more optimized misprediction mitigation, LAVA achieves best average improvement of 1.5 pp compared to LA-Binary.
We also observe that LAS consistently outperforms LA when both have oracular lifetimes: 9.5 pp vs. 7.5 pp on average, confirming our theoretical results that reprediction is fundamentally more powerful than one-shot prediction.
6.2 Production Results (LAS)
Prior to our fleet-wide roll-out, we ran several waves of production pilots on C2 and E2. For the first two waves, we conducted A/B experiments in three pools and measure the impact. We observe statistically significant empty host improvements ranging from 2% to 10% for the entire pool, shown in Table 1.
Table 1: Empty Host Improvements in Pilot Pools
Pilot Pool Type Change in Empty Hosts
C2 Wave 1 pool A/B +2.3 pp (p-value = 0.01)
C2 Wave 2 pool 1 A/B +2.7 pp (p-value $<$ 0.01)
C2 Wave 2 pool 2 A/B +9.2 pp (p-value $<$ 0.01)
C2 Wave 3 pool All +4.9 pp (95% CI: [0.54, 9.2])
E2 Wave 1 pool All +6.1 pp (95% CI: [1.9, 10.0])
Table 2: VM Migration Reductions Using LARS on Two Traces.
Migrations
Scheduled Baseline LARS Reduction
1 48,239 37,108 35,505 4.32%
2 53,597 36,307 34,655 4.55%
For the third wave, we piloted the entire pool with LAS. We employ a Bayesian Structural Time Series method CausalImpact (Brodersen et al., 2015) to evaluate the pre-post causal effect brought by LAS (Figure 7). We observed a significant empty host improvement of 4.9% (Table 1). Moreover for the whole-pool pilot, we additionally measure the causal effect of LAS on stranded CPU and memory, where 1 pp of stranding reduction translates directly into 1% of capacity. LAS reduced CPU stranding by 3% (95% CI: [-5.5%, -0.49%]) and memory stranding by 2% (95% CI: [-3.8%, -0.23%]). We also conducted a whole-pool pilot for an E2 pool.
6.3 Lifetime-based Host Defragmentation (LARS)
Using production traces, we simulate defragmentation on two different one-month time intervals from a number of pools with customer workloads, including C2 and E2 VMs. In this experiment, we use oracle lifetimes. LARS does not affect which hosts are picked for migration at a given time, but orders the VMs across all in-migration hosts so that the VMs with the longest remaining lifetime are migrated first. The effect is that a larger number of VMs terminate before they would have been migrated, resulting in a reduction of around 4.5% of live migrations.
6.4 ML Model
(a) Empty hosts.
(b) CPU stranding.
Figure 7: Time series from our Wave 3 rollout, produced by the CausalImpact (Brodersen et al., 2015) package. Each sub-figure has three panels. The first panel shows the data and a counterfactual prediction (dotted line) for the post-launch period. The second panel shows the difference between observed data and counterfactual predictions, i.e., the point-wise causal effect, as estimated by the model. The third panel adds the point-wise contributions from the second panel, plotting the cumulative effect.
Our model achieves 70% recall at 99% precision, which is significantly higher than other model types we tried. We measured the latency of our model and find that most predictions complete in under 10 us (Figure 8). This latency improves upon Barbalho et al. (2023) by 780$\times$, and shows the benefit of compiling the model into the Borg Prime binary. Fast predictions make it practical to frequently re-run the model to correct mispredictions and to use it during maintenance and defragmentation.
6.5 Impact of Repredictions
To further understand the utility of repredictions leveraging the current VM uptime, we vary the amount of uptime used as input feature to the model and assess its impact on model accuracy. For each VM in the test trace, we divide its total lifetime into 20 quantiles. We then show the F1 score for predicting whether or not a VM lives beyond 168 hours (7 days) when the uptime is set to a particular quantile. Figure 9 illustrates the relationship between varying amount of uptime and model accuracy. We observe that without using reprediction (no uptime/0th quantile), the model only has a F1 score of 0.8. The F1 score quickly rises above 0.9 after the 8th quantile (i.e., once the VM’s uptime has reached 40% of its total lifetime). Observing how long a VM has lived, and repredicting its remaining lifetime thus leads to more accurate predictions. We find repredicting longer-lived VMs is critical to avoid spreading them across hosts and causing fragmentation.
Figure 8: Histogram of model execution latencies.
Figure 9: Making repredictions of VM remaining lifetime using its uptime significantly improves the model accuracy.
Figure 10: The model remains high accuracy weeks after training, making its maintenance overhead low for production deployment.
7 Related Work
Prediction-based job scheduling has a long history. Some work uses RL Mao et al. (2019); Li et al. (2021); Wang et al. (2022) to learn decisions directly, such as for placing containers. Other work predicts properties of jobs and uses them to improve scheduling for scale-out Delimitrou & Kozyrakis (2014a), interference Delimitrou & Kozyrakis (2013), or goodput Jayaram Subramanya et al. (2023).
Job lifetimes are particularly useful for scheduling. Jockey Ferguson et al. (2012) and Aria Verma et al. (2011) predict runtimes in data processing frameworks and use them to improve scheduling of dataflow-based jobs. TetriSched Tumanov et al. (2016) uses runtime predictions in conjunction with an MILP solver to schedule jobs in an HPC-style setting. Wrangler uses them to identify stragglers Yadwadkar et al. (2014). 3Sigma Park et al. (2018) predicts various properties of Google job traces and uses them for scheduling. Note that most of this work uses runtime predictions in domain-specific scenarios and does not leverage general machine learning techniques to schedule.
In Cloud Computing, Microsoft’s Resource Central system Cortez et al. (2017) performs predictions on VMs and uses them for various tasks, including VM scheduling and admission control Sajal et al. (2023). Microsoft’s LA algorithm that performs on lifetime-based VM scheduling is the most similar to ours Barbalho et al. (2023); Buchbinder et al. (2022). Our work differs in that instead of performing one-shot predictions when the VM starts running, it repredicts VM lifetimes as needed, updating lifetimes based on new knowledge, delivering good improvements in empty hosts over LA. We show how to incorporate it in our existing scheduler by integrating the models into the binary. Our models run in-process at very low latency and our new scheduling algorithms tolerate mispredictions.
8 Conclusion
We presented a new approach to lifetime-aware VM allocation in cloud data centers. Our approach continually repredicts the lifetime of VMs and corrects for past mistakes. LAS is deployed fleet-wide for one VM family at Google, where it shows significant improvements both in terms of increasing empty hosts and reducing stranding.